# Policies for multiplexed quantum repeaters: theory and practical performance analysis

### v1

@@ Abstract
Future quantum networks will have nodes equipped with multiple quantum memories, providing the possibility to perform multiplexing and distillation strategies in order to increase fidelities and reduce waiting times for end-to-end entanglement distribution. In this paper, we introduce two policies that adapt the well-known swap-as-soon-as-possible (swap-asap) policy to multiplexed quantum repeater chains. Unlike the usual, fully local swap-asap policy, these policies are “quasi-local”, making effective use of knowledge of the states of the repeaters along the chain to optimize waiting times and end-to-end fidelities. Our policies also make use of entanglement distillation. We demonstrate via simulations one of our key findings, which is that these policies can outperform the well-known and widely studied nested purification and doubling swapping policy in practically relevant parameter regimes. Our work also provides the tools to carefully examine the role of entanglement distillation. We identify the parameter regimes in which performing distillation makes sense and is useful. In these regimes, we also address the question: “Should we distill before swapping, or vice versa?” We thus formalize the trade-off between the advantages of adding distillation capabilities to quantum networks against their technological and practical challenges. Finally, to provide further practical guidance, we propose an experimental implementation of a multiplexing-based linear network, and experimentally demonstrate the key element, a high-dimensional biphoton frequency comb (BFC). We then evaluate the anticipated performance of our multiplexing-based policies in such a real-world network through simulation results for two concrete memory platforms, namely rare-earth ions and diamond vacancies.
@@ I Introduction
The quantum internet [1, 2, 3] has the potential to revolutionize current computation, communication, and sensing technologies by enabling exchange of quantum data. Numerous significant applications have already been recognized, encompassing areas such as secure quantum communication [4, 5], distributed quantum computation [6, 7], and distributed quantum sensing [8, 9, 10], to name a few. However, enormous hardware improvements are needed before a practical quantum internet is realized [11, 12]. Currently, a significant challenge lies in effectively distributing entanglement amongst network nodes, particularly over significant distances. Recent, state-of-the-art experiments have been performed for only a handful of nodes over short distances [13, 14]. This challenge stems from the delicate nature of quantum systems, leading to issues like photon losses, imperfect measurements, and quantum memories with short coherence times. Since comprehensive, fault-tolerant quantum error correction has yet to be realized to overcome these obstacles, it becomes important to explore how much we can overcome these limitations through other means. This problem of designing effective protocols for entanglement distribution using noisy, imperfect quantum hardware has been addressed in multiple analytical [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and numerical [34, 35, 36, 37, 38, 39, 40, 41, 42, 43] studies, among others; see also Refs. [44, 45, 46, 47, 48, 49, 50] for reviews.
In the context of entanglement distribution in quantum repeater chains, the “single-channel” approach has been commonly studied. In the single-channel setting, every repeater is connected via a single channel to each of its two neighbors, with a single quantum memory at the endpoint of each channel. These channels are used to create elementary links of entangled qubit pairs between the neighboring nodes. Then, using a particular policy for performing probabilistic entanglement-swapping operations [51, 52], the elementary links are transformed into a virtual link between the two end nodes of the repeater chain, which can be further used for exchanging quantum information via the quantum teleportation protocol [51]. Noise and loss along the channels, along with quantum memories with short coherence times, result in long waiting times and/or low fidelities of the end-to-end entangled links.
@@
Figure 1: (a) Difference between the parallel and multiplexing approaches. Unlike the parallel approach, cross-channel elementary link generation and entanglement swapping is permitted in the multiplexing approach. (b) Elementary links are created by multiple entanglement generation sources associated with every channel that connects different pairs of quantum memories. (c) Non-neighbouring nodes are connected by virtual links generated by performing entanglement swapping operations between neighbouring active quantum memories. (d) Linear chain of quantum repeaters with multiple quantum memories capable of establishing multiple physical links, entanglement distillation and entanglement swapping.
In this work, we look at the multiplexing approach [17] as a potential method for reducing the overall waiting time for entanglement generation in linear networks and increasing the fidelity of end-to-end entangled links; see Fig. 1. In a multiplexing protocol, every repeater contains multiple memories, which are connected by multiple channels, as shown in Fig. 1(a). This allows for simultaneous elementary link generation attempts. Furthermore, when performing entanglement swapping, it is possible to go beyond the trivial parallel approach and perform cross-channel entanglement swapping; see Fig. 1(a). There is also the possibility to perform entanglement purification/distillation [53, 54, 55, 56].
Despite the significant benefits of approaches based on multiplexing and/or distillation already studied and demonstrated in the literature [15, 16, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74], many open questions remain in terms of efficient and optimal entanglement swapping and distillation policies for linear quantum repeater chains. In particular, many of the aforementioned works on multi-memory quantum repeaters with multiplexing have made use of the “nested purification and doubling swapping” protocol devised in the seminal works [15, 16] on quantum repeaters. We therefore ask the following two questions.
@@
    What is the best way to combine entanglement swapping with entanglement distillation? Should we distill first before swapping, or vice versa?
I.1 Overview of results
In this work, we start by extending the framework from Ref. [43] for modeling quantum repeater chains in terms of Markov decision processes to the case of multi-memory quantum repeater chains (see Sec. II). Then, we use this model to address the two questions above.
Our main results and contributions comprise four parts.
  1. 1.
    New policies for multiplexed quantum repeaters.
    In Sec. III, we define our multiplexing-based policies. These policies are generalizations of the so-called “swap-as-soon-as-possible” (swap-asap) policy [29, 30] to the multiplexing setting. We refer to them as strongest neighbor (SN) swap-asap, farthest neighbor (FN) swap-asap, and random swap-asap. These policies are distinguished by how they pair memories within nodes for the purposes of entanglement swapping. Intuitively, the SN swap-asap policy prioritizes the creation of high-fidelity end-to-end links, and FN swap-asap prioritizes lowering the waiting time for the creation of end-to-end links. These policies are also “quasi-local”, requiring some knowledge of the network state but not full, global knowledge.
    In Sec. IV, we compare the performance of these policies and show that the SN, FN, and random swap-asap policies can significantly reduce the average waiting time compared to the parallel, non-multiplexed version of swap-asap, and at the same time provide higher end-to-end link fidelities.
    Notably, the SN and FN policies can outperform the widely used doubling protocol for entanglement swapping in certain parameter regimes. More specifically, in Sec. IV.1, we see improved performance of our policies in more resource-constrained scenarios, such as high channel losses, high qubit noise (small coherence time of quantum memories), and when the number of multiplexing channels is small and/or the number of network nodes is large. These parameter regimes are relevant for near-term quantum network implementations.
  2. 2.
@@ I.1 Overview of results
    Then, in Sec. V, we add distillation to our FN and SN swap-asap policies, and we address the following questions: i) “When is distillation useful?” ii) “Should we distill before swapping, or swap first and then distill?” We demonstrate via simulations that these policy considerations have a strong influence on the performance of the network. As such, with our study, we provide the tools to weigh the technological and practical challenges of adding distillation abilities to quantum networks against the advantages of adding such an ability.
  3. 3.
    The impact of classical communication costs.
    In Sec. VI, we consider the classical communication (CC) costs associated with our policies. By incorporating the time needed for CC into our model, we show that our multiplexing-based FN and SN swap-asap policies outperform fully-local policies that do not require any CC. We thus show that the benefit gained by acquiring information about the network’s state is retained even when the communication costs of acquiring such information is accounted for. This result provides much needed evidence in support of exploring policies that use knowledge of the network state, as promoted in previous literature [42, 41, 43]. We also show that, even when CC costs are accounted for, the FN and SN swap-asap policies can outperform the doubling policy in parameter regimes corresponding to resource-constrained hardware settings.
@@
    In Sec. VII, we outline a proof-of-principle experimental realization of a quantum network equipped with multiplexing policies using a high-dimensional bi-photon frequency comb (BFC), and we map the relevant experimental parameters to their counterpart parameters within our modeling framework. Then, in Sec. VII.2, we assess the performance of our multiplexing-based policies within these parameter regimes for two quantum memory platforms, diamond vacancy and rare-earth metal ions.
@@ II Linear chain quantum network with multiple channels
  • Elementary and virtual links: Entanglement sources establish physical/elementary links between nearest-neighbour quantum memories with probability $p_{\ell}\in[0,1]$. A probabilistic Bell state measurement creates a virtual link between distant quantum memories with success probability $p_{sw}\in[0,1]$. These values are fundamentally limited by the technology used to create the system, e.g., linear optics makes $p_{sw}\leq 0.5$, and $p_{\ell}$ is determined by properties of the source, loss characteristics of the sources, etc. These considerations are relevant for our proposed experimental implementation in Sec. VII, and we provide details on how to determine $p_{\ell}$ and $m^{\star}$ from physical properties of the system in Sec. VII.1.
    We keep track of the fidelity of both elementary and virtual links by assigning every link an age. The age $m$ of a link starts from $m=m_{0}$, when it is first created, with $m_{0}$ being a measure of its initial fidelity with respect to a perfect Bell state. The age then increases in discrete steps, such that $m\in\{0,1,2,3...\}$. Once the age of the link reaches $m^{\star}$, the link is discarded. Throughout this work, we consider a particular Pauli noise model for decoherence of the memories (see Appendix A), such that the fidelity $f(m)$ of a link that has age $m$ is given by
@@
    The age of a link formed by a successful entanglement swapping operation is determined by adding the ages of the two corresponding links: if $m_{1}$ and $m_{2}$ are the ages of the two links, then if the entanglement swapping is successful the age of the new link is $m^{\prime}=m_{1}+m_{2}$. This “addition rule” for the age of virtual links is a direct consequence of the Pauli noise model that we use, and a proof can be found in Ref. [43, Appendix C].
  • Entanglement distillation: Multiple low-fidelity links between any two nodes can be distilled with probability $p_{ds}\in[0,1]$ to form a higher-fidelity link. Throughout this work, we consider the BBPSSW protocol [53], and we provide details on it in Appendix B.
Figure 2: (a) Illustrative example of the FN swap-asap policy. Links on both sides of the $k^{\text{th}}$ node are ranked separately on the basis of their lengths. The longest link is given a rank of 1, while shortest is ranked 3. For entanglement swapping, we thus give first preference to links of rank 1 (marked in red), which would result in the longest possible link. Being a probabilistic process, if the entanglement swapping fails, then next preferences is given to links of rank 2, and so on. (b) In the case of the SN swap-asap policy, link rankings are based on their ages. Preference is given to links with lower ages.
Entanglement distribution protocols progress in a series of discrete time steps, based on the Markov decision process model developed and used in Ref. [43], which we refer to for further details on the model. In every time step, the following events occur.
@@
    If entanglement distillation is part of the policy, attempt to distill all links between the same two nodes for a given number of rounds, as dictated by the policy; see Sec. V for details. If the policy calls for distilling first, then swapping, interchange steps 2 and 3.
@@
    Increase the age of all active links by one time unit, accounting for decoherence. If classical communication (CC) overheads are to be accounted for, then add the CC time to the age of every active link; see Sec. VI for details. Discard any link which has age greater than $m^{\star}$
III Policies for multiplexing
In this section, we introduce our two multiplexing policies for entanglement distribution in a linear chain quantum network.

### v2

# Reducing classical communication costs in multiplexed quantum repeaters using hardware-aware quasi-local policies

@@ Abstract
Future quantum networks will have nodes equipped with multiple quantum memories, allowing for multiplexing and entanglement distillation strategies in order to increase fidelities and reduce waiting times for end-to-end entanglement distribution. In this work, we introduce quasi-local policies for multiplexed quantum repeater chains. In fully-local policies, nodes make decisions based only on knowledge of their own states. In our quasi-local policies, nodes have increased knowledge of the state of the repeater chain, but not necessarily full, global knowledge. Our policies exploit the observation that for most decisions the nodes have to make, they only need to have information about the connected region of the chain they belong to, and not the entire chain. In this way, we not only obtain improved performance over local policies, but we reduce the classical communication (CC) costs inherent to global-knowledge policies. Our policies also outperform the well-known and widely studied nested purification and doubling swapping policy in practically relevant parameter regimes. We also carefully examine the role of entanglement distillation. Via analytical and numerical results, we identify the parameter regimes in which distillation makes sense and is useful. In these regimes, we also address the question: “Should we distill before swapping, or vice versa?” Finally, to provide further practical guidance, we propose an experimental implementation of a multiplexing-based repeater chain, and experimentally demonstrate the key element, a high-dimensional biphoton frequency comb. We then evaluate the anticipated performance of our multiplexing-based policies in such a real-world network through simulation results for two concrete memory platforms, namely rare-earth ions and diamond vacancies.
@@ I Introduction
The quantum internet [1, 2, 3] has the potential to revolutionize current computation, communication, and sensing technologies by enabling exchange of quantum data. Numerous significant applications have already been recognized, encompassing areas such as secure quantum communication [4, 5, 6], distributed quantum computation [7, 8], and distributed quantum sensing [9, 10, 11], to name a few. However, enormous hardware improvements are needed before a practical quantum internet is realized [12, 13]. Currently, a significant challenge lies in effectively distributing entanglement amongst network nodes using quantum repeaters, with the goal of attaining high fidelities and low waiting times, particularly over significant distances. Recent, state-of-the-art experiments have been performed for only a handful of nodes over short distances [14, 15]. This challenge stems from the delicate nature of quantum systems, leading to issues like photon losses, imperfect measurements, and quantum memories with short coherence times. Since comprehensive, fault-tolerant quantum error correction has yet to be realized to overcome these obstacles, it becomes important to explore how much we can overcome these limitations through other means. This problem of designing effective repeater protocols for entanglement distribution using noisy, imperfect quantum hardware has been addressed in multiple analytical [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] and numerical [36, 37, 38, 39, 40, 41, 42, 43, 44, 45] studies, among others; see also Refs. [46, 47, 48, 49, 50, 51, 52] for reviews.
In order to make the best use of today’s noisy, imperfect quantum hardware, protocols/policies for entanglement distribution should be hardware-aware. This means tailoring the elementary link generation and entanglement swapping policies to the hardware, instead of employing widely-studied but hardware-agnostic policies such as the doubling [16, 17] and the “swap-as-soon-as-possible” [30, 31] policies  [^1] Another consideration is tailoring the number and locations of repeater stations and entanglement sources to the available hardware, which are architectural questions that have have been deeply studied; see, e.g., Refs. [84, 62, 3, 59, 100, 101].. Determining optimal hardware-aware policies for elementary link generation and entanglement swapping has received heightened attention only recently [54, 29, 55, 56, 43, 44, 45, 57].
At the same time, for the first-generation quantum repeaters that we consider here, it is well known that the two-way end-to-end classical communication involved makes them rather slow as the end-to-end distance increases, when compared to second- and third-generation repeaters, which use quantum error-correction [58, 59]. Nevertheless, first-generation quantum repeaters are arguably the most amenable to implementation in the near future. Therefore, it is of interest to develop policies for elementary link generation and entanglement swapping that are not only hardware-aware, but also minimize the amount of classical communication between nodes in the network.
The amount of classical communication in a network is tied to the amount of knowledge nodes have of each other’s states when making their decisions regarding elementary link generation and entanglement swapping. Indeed, the more knowledge nodes need to make use of, the more classical communication is required to attain that knowledge. The decisions that nodes make, and the amount of knowledge needed to make the decisions, is dictated by the policy being used. In particular, if a policy calls for the nodes to make use of full, global knowledge of the network—meaning that every node should know the state of every other node at all times—then classical communication over the entire span of the network is required in every time step of the protocol. On the other extreme, for fully local policies, in which nodes need to have knowledge only of their own states, only one round of end-to-end classical communication is required at some prescribed final time. In the middle lie quasi-local policies, which we introduce in this work. These policies use to their advantage the realization that nodes only need access to knowledge of the connected portion of the chain they are part of. Disconnected regions of the network can determine their actions independently. Thus, barring only a few time steps when long links are present in the network, spanning lengths on the order of the size of the network, full end-to-end classical communication is superfluous and wasteful. To the best of our knowledge, a thorough analysis of end-to-end waiting times and fidelities, as a function of the amount of knowledge nodes have of each other’s states, has not been conducted. Recent works [44, 45] have highlighted the importance of global knowledge in improving waiting times and fidelities, but without fully accounting for the cost of classical communication. We are therefore motivated to ask our first question:
  1. (Q1)
    How are waiting times and fidelities affected by the amount of classical communication being performed in a network? Are the benefits of additional network knowledge, beyond that of a fully local policy, negated by the increase in classical communication?
In this work, we consider this question in the context of entanglement distribution in linear networks (also called “repeater chains”) with multi-memory quantum repeaters and multiplexing [16, 17, 60, 18, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]; see Fig. 1. In the multiplexing approach, simultaneous elementary link generation attempts can be made, and when performing entanglement swapping [5, 78], it is possible to go beyond the trivial parallel approach and perform cross-channel entanglement swapping.
Many of the aforementioned works on multi-memory quantum repeaters with multiplexing have made use of the hardware-agnostic “nested purification and doubling swapping” protocol devised in the seminal works [16, 17] on quantum repeaters. Keeping in mind our goal of developing hardware-aware policiesg, we therefore ask:
  1. (Q2)
In the multiplexing approach, there is also the possibility to perform entanglement purification/distillation [79, 80, 81, 82]. Distillation introduces addition classical communication overhead. We are therefore motivated to ponder whether the potential benefits of distillation on end-to-end fidelities are worth the additional classical communication overhead. Specifically, we ask:
  1. (Q3)
    In what parameter regimes is it beneficial to perform entanglement distillation? What is the best way to combine entanglement swapping with entanglement distillation—should we distill first before swapping, or vice versa?
@@
Figure 1: (a) In the multiplexing approach to entanglement distribution, it is possible to perform cross-channel elementary link generation and entanglement swapping. (b) Elementary links are created by multiple entanglement generation sources associated with every channel that connects different pairs of quantum memories. (c) Non-neighbouring nodes are connected by virtual links generated by performing entanglement swapping operations between neighbouring active quantum memories. (d) A linear chain of quantum repeaters with multiple quantum memories capable of establishing multiple physical links, entanglement distillation and entanglement swapping.
I.1 Overview of results
We address the three questions above by extending the framework from Ref. [45] for modeling quantum repeater chains in terms of Markov decision processes to the case of multi-memory quantum repeater chains (see Sec. II). Our main results and contributions comprise four parts.
@@ I.1 Overview of results
    Quasi-local policies for multiplexed quantum repeaters.
    In Sec. III, we define our quasi-local multiplexing-based policies. These policies are generalizations of the so-called “swap-as-soon-as-possible” (swap-asap) policy [30, 31] to the multiplexing setting. We refer to them as strongest neighbor (SN) swap-asap, farthest neighbor (FN) swap-asap, and random swap-asap. These policies are distinguished by how they pair memories within nodes for the purposes of entanglement swapping. Intuitively, the SN swap-asap policy prioritizes the creation of high-fidelity end-to-end links, and FN swap-asap prioritizes lowering the waiting time for the creation of end-to-end links. These policies are also “quasi-local”, requiring some knowledge of the network state but not full, global knowledge.
@@
    The impact of classical communication costs.
    In a realistic entanglement distribution setting with imperfect memories, considering classical communication (CC) costs is imperative in order to provide a realistic performance analysis. We incorporate the CC costs associated with our policies, and thereby address (Q1). In particular, we show in Sec. IV that our multiplexing-based FN and SN swap-asap policies outperform fully-local policies that do not require any CC. We thus show that the benefit gained by acquiring information about the network’s state is retained even when the communication costs of acquiring such information is accounted for.
    We also address (Q2) by showing that the FN and SN swap-asap policies can outperform the doubling policy in parameter regimes corresponding to high channel losses, small coherence times of quantum memories, and a small number of multiplexing channels and/or a large number of nodes.
  3. 3.
@@
    We also add entanglement distillation to our FN and SN swap-asap policies, and in Sec. V we address (Q3). We find that if the fresh elementary links have high fidelity, using entanglement distillation turns out to be detrimental. On the other hand, when fresh elementary links have low fidelity, and more generally in resource-constrained scenarios, entanglement distillation can be more useful. Nonetheless, even in such cases, we observe examples in which distillation can in fact be detrimental.
    In terms of the relative ordering of distillation and swapping, if links are have low generation probability or low initial fidelity, then distilling first is more advantageous. When the elementary link generation probability is high, it is better to swap first and then distill.
@@
    In Sec. VI, we outline a proof-of-principle experimental realization of a quantum network equipped with multiplexing policies using a high-dimensional bi-photon frequency comb (BFC), and we map the relevant experimental parameters to their counterpart parameters within our modeling framework. Then, in Sec. VI.2, we assess the performance of our multiplexing-based policies within these parameter regimes for two quantum memory platforms, diamond vacancy and rare-earth metal ions. In particular, we consider the number of repeater nodes needed to achieve a desired end-to-end waiting time and fidelity over a distance of 100 km.
@@ II Linear chain quantum network with multiple channels
  • Elementary and virtual links: Entanglement sources establish physical/elementary links between nearest-neighbour quantum memories with probability $p_{\ell}\in[0,1]$. A probabilistic Bell state measurement creates a virtual link between distant quantum memories with success probability $p_{sw}\in[0,1]$. These values are fundamentally limited by the technology used to create the system, e.g., linear optics makes $p_{sw}\leq 0.5$, and $p_{\ell}$ is determined by properties of the source, loss characteristics of the sources, etc. These considerations are relevant for our proposed experimental implementation in Sec. VI, and we provide details on how to determine $p_{\ell}$ and $m^{\star}$ from physical properties of the system in Sec. VI.1.
    We keep track of the fidelity of both elementary and virtual links by assigning every link an age. The age $m$ of a link starts from $m=m_{0}$, when it is first created, with $m_{0}$ being a measure of its initial fidelity with respect to a perfect Bell state. The age then increases in discrete steps, such that $m\in\{0,1,2,3...\}$. Once the age of the link reaches $m^{\star}$, the link is discarded. Throughout this work, we consider a particular Pauli noise model for decoherence of the memories (see Appendix B), such that the fidelity $f(m)$ of a link that has age $m$ is given by
@@
    The age of a link formed by a successful entanglement swapping operation is determined by adding the ages of the two corresponding links: if $m_{1}$ and $m_{2}$ are the ages of the two links, then if the entanglement swapping is successful the age of the new link is $m^{\prime}=m_{1}+m_{2}$. This “addition rule” for the age of virtual links is a direct consequence of the Pauli noise model that we use, and a proof can be found in Ref. [45, Appendix C].
  • Entanglement distillation: Multiple low-fidelity links between any two nodes can be distilled with probability $p_{ds}\in[0,1]$ to form a higher-fidelity link. Throughout this work, we consider the BBPSSW protocol [79], which allows for distillation of two entangled links to one. We provide a summary of the protocol in Appendix C. Briefly, if $f_{1}$ and $f_{2}$ are the fidelities of the two links being distilled, then the success probability $P_{\text{distill}}(f_{1},f_{2})$ of the protocol and the resulting fidelity $F_{\text{distill}}(f_{1},f_{2})$ are given by
    We show in Appendix C that the age after distillation with the BBPSSW protocol is given by the following formula:
    $m^{\prime}=\left\lceil m^{\star}\log\left(\frac{15-6(f_{1}+f_{2})+24f_{1}f_{2}}{32f_{1}f_{2}-2(f_{1}+f_{2})-1}\right)\right\rceil,$ (4)
    where $f_{1}\equiv f(m_{1})$, $f_{2}\equiv f(m_{2})$, $m_{1}$ and $m_{2}$ are the ages of the two links being distilled, the function $m\mapsto f(m)$ is defined in Eq. (1), and $\lceil x\rceil$ is the smallest integer greater than or equal to $x$.
Entanglement distribution protocols progress in a series of discrete time steps, based on the Markov decision process model developed and used in Ref. [45], which we refer the reader to for further details on the model. In every time step, the following events occur.
@@
    If entanglement distillation is part of the policy, attempt to distill all links between the same two nodes until either one or no link survives; see Sec. III for details. If the policy calls for distilling first, then swapping, interchange steps 2 and 3.
@@
    Increase the age of all active links by one time unit, accounting for decoherence. If classical communication (CC) overheads are to be accounted for, then add the CC time to the age of every active link; see below for details. Discard any link which has age greater than $m^{\star}$.
The time steps that we consider can be converted into physical times, and also the ages can be converted to fidelity values, by incorporating the network hardware and design parameters. Examples of such a translation are given in Sec. VI.1 and further details are provided in Appendix B.
II.1 Classical communication
Since entanglement distribution at the elementary link level is not deterministic, it must be heralded via classical signals. This heralding sets a natural time scale for entanglement distribution through a network. The heralding time is equal to the classical communication time between two adjacent nodes. For the purposes of our simulation, this heralding time becomes the duration of one time step. Since entanglement swapping is also often probabilistic, and entanglement distillation requires two-way classical communication [80] between nodes, two distinct classes of entanglement distribution approaches can be envisaged.
  • A local approach, where all elementary link generation and entanglement swapping attempts are made agnostic to the success or failure of prior swaps. This would then mean that all elementary links (once generated successfully) are retained up to the cutoff time. When an entanglement swap has failed, resulting in both participating links to become inactive, the surrounding nodes must act as if the link is still active until it has reached its expected cutoff. All the classical communication about successful swaps is then done at the end of the protocol. See Appendix F for details about how the end of the protocol is established. Such an approach is fully local and therefore all policy decisions are taken by individual nodes, without any collaboration or exchange of information. The benefit is that no classical communication overheads exist except for heralding of elementary links, and therefore every time step has duration equal to the heralding time only. The cost is instead paid by having to wait for a link to reach its cutoff time, even though it might have become inactive long before that. Distillation can be only included for freshly generated elementary links in a fully local scheme, since as links get older their lengths cannot be predicted locally (since results of swaps are not known).
In previous works such as Refs. [44, 45], global policies were shown to optimize the average waiting time and fidelity; however, classical communication costs were not included. Furthermore, in Ref. [45], a “quasi-local” approach to entanglement distribution policies was proposed, in which there is knowledge of the network state up to some length scales only, and not globally. In this case, an advantage over fully local policies could still be obtained. We now investigate in this work whether the advantages in waiting time and fidelity gained by global and quasi-local knowledge and collaboration can be retained when CC overheads are taken into account.
Below is a summary of how we have taken classical communication into account in our model  [^2] Let us make a brief remark. In a full simulation, such as those which can be done via quantum network simulators such as NetSquid [37], QuISP [40], SeQUeNCe [38], etc., time stamps are assigned to each quantum operation and tasks are queued, such that classical communication costs are intrinsically kept track of. Our simulations focus on optimization of policies, similar in spirit to rule sets used by many of the above mentioned simulators; see also Refs. [102, 103]. .
  • We approximate the CC time by adding $t_{cc}$ time steps to every network evolution step, where $t_{cc}$ is equal to the length of the longest link (number of nodes) involved in any entanglement swap in that step times the CC time between two adjacent nodes of the chain. This allows enough time for the classical communication for all entanglement swaps attempted in that MDP step, and at the end of such a CC-accounted MDP step, all nodes are aware of which nodes they are connected to and the ages of those links (which can also be classically communicated). In most cases this estimate acts as an upper bound to the CC cost, since in practice nodes could perform the next required action as soon as they receive the CC tagged for or relevant to them. See Appendix F for details.
  • CC costs associated with entanglement distillation are added based on the length of the longest links that are distilled in a time step. For local policies, distillation is only allowed for freshly prepared elementary links (if they are not perfect), and thus the CC cost is always equal to one time step, if distillation is performed, between any two adjacent nodes after heralded entanglement generation.
Figure 2: (a) Illustrative example of the FN swap-asap policy. Links on both sides of the $k^{\text{th}}$ node are ranked separately on the basis of their lengths. The longest link is given a rank of 1, while shortest is ranked 3. For entanglement swapping, we thus give first preference to links of rank 1, which would result in the longest possible link. If entanglement swapping fails, then next preference is given to links of rank 2, and so on. (b) In the case of the SN swap-asap policy, link rankings are based on their ages. Preference is given to links with lower ages.
III Quasi-local multiplexing policies